OpenStudy (anonymous):

Solve the system of equations. 4x + 7y = 6 and 6x + 5y = 20

OpenStudy (jdoe0001):

have you covered "solving by elimination" yet?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 4x + 7y = 6& {\color{red}{ \times -5}}\implies &-20x\cancel{-35y}=-30\\ 6x + 5y = 20& {\color{red}{ \times 7}}\implies &\quad 42x\cancel{+35y}=140\\ \hline\\ &&\quad \square x\qquad \qquad =\square \end{array}\) solve for "x" then substitute that found "x" on either equation to get "y"